# Controlling Large Language Model-based Agents for Large-Scale Decision-Making: An Actor-Critic Approach

### 23 Nov 2023 v1

@@ 1 Introduction
Relying on training from massive datasets to capture extensive common knowledge and having demonstrated certain reasoning capabilities, Large Language Models (LLMs) have been widely applied and explored across various domains, rapidly emerging as powerful tools Brown et al. (2020); Kojima et al. (2022); Ruan et al. (2023a); Yang et al. (2023). Examples of their applications include question-answering systems Mallen et al. (2023), common-sense reasoning Hao et al. (2023), programming Tian et al. (2023), and embodied intelligence Driess et al. (2023). Recently, in the fields of natural language processing (NLP) and multi-agent systems (MAS), there has been a growing interest in exploring the use of multiple LLMs to collaboratively accomplish tasks. However, the majority of existing works primarily concentrate on coordination among a limited number of agents. The application of LLMs for effective coordination in large-scale agent scenarios has been rarely discussed.
@@
To this end, we present LLaMAC, a novel framework for achieving a comprehensive decision-making process in collaborative tasks involving large-scale LLM-based agents, drawing inspiration from the classical actor-critic reinforcement learning (RL) approach Konda and Tsitsiklis (1999). Within this framework, a centralized critic takes on the role of a coordinator, making suggestions to each actor based on their decision memory. Subsequently, the actors interact with the environment, receive assigned tasks, and execute corresponding actions. (1) To tackle the exploration-exploitation trade-off inherent in the decision-making process, we introduce a TripletCritic structure inspired by the distributional code for value in the brain Dabney et al. (2020). This structure coordinates multiple critics with the same objective but different preferences. (2) We also establish a comprehensive feedback mechanism that incorporates both internal feedback within the TripletCritic and external feedback between the LLM-based actors (i.e., agents) and the TripletCritic. This mechanism aims to mitigate hallucination issues and bolster the robustness of LLM decision-making. (3) To construct a token-efficient planning framework, we incorporate mechanisms such as redundant memory information deletion and external feedback judgment. These mechanisms effectively reduce the length of context tokens and minimize the frequency of LLM accesses.
We first evaluate the performance of our method on a system resource allocation task to demonstrate its ability to strike a balance between exploration and exploitation, as well as its capability in large-scale multi-agent decision-making tasks. Subsequently, we deploy our method in a more complex robot grid transportation scenario to validate its planning and decision-making capabilities. Experimental results demonstrate that our method outperforms existing approaches in terms of final performance, token utilization efficiency, and policy stability. To the best of our knowledge, we are the first to apply LLM to large-scale multi-agent decision-making tasks.
@@ 2 Related Work
2.1 Multi-Agent Cooperation.
@@
2.2 Planning and Reasoning with LLM.
Learning in massive corpora gives LLMs certain commonsense reasoning capabilities Kojima et al. (2022). Although there are still challenges in solving complex decision tasks, a large amount of work has proven that their methods can effectively improve the planning ability of LLMs Zelikman et al. (2022); Creswell et al. (2022). One line of research focuses on decomposing complex queries into sequential intermediate steps, known as Chain-of-Thought (CoT) Wei et al. (2022), to achieve accurate solutions. Another direction involves incorporating feedback mechanisms, showcasing their extensive capabilities in tackling complex decision-making challenges Wang et al. (2023). Moreover, recent studies have begun to address this issue employing multiple LLMs. These approached are enhanced in their planning capabilities through techniques such as debate Chan et al. (2023); Liang et al. (2023) or role-playing Li et al. (2023); Hong et al. (2023). In the domain of decision-making, a subset of research utilizes prompting techniques to construct comprehensive processes covering perception, planning, and action, in cluding video games Zhang et al. (2023c), robot contro l Zhang et al. (2023d); Mandi et al. (2023), and open-world tasks Zhu et al. (2023); Gong et al. (2023). There are some studies about task planning and external tool usage Ruan et al. (2023a, b); Kong et al. (2023). Our work specifically focuses on the application of language models in large-scale Multi-Agent System decision-making.
@@ 4.2 Grid Transportation
Success Steps Feedback Token
HMAS-2 100% 9.9(2.74) 3.3(2.05) 49877.7(17984.51)
2x2 LLaMAC 100% 7.0(1.79) 2.0(1.26) 23921.2(8382.90)
HMAS-2 80% 15.5(6.09) 12.3(5.83) 158385.9(107838.06)
2x4 LLaMAC 100% 7.6(1.36) 4.3(1.42) 37984.9(10570.54)
HMAS-2 60% 30.6(9.70) 26.1(13.59) 599252.9(245398.21)
4x8 LLaMAC 100% 12.9(2.70) 10.7(3.35) 122607.0(30545.11)
Success Steps Feedback Token
HMAS-2 80% 7.0(5.0) 6.0(9.74) 76140.4(116655.08)
2x2 LLaMAC 100% 4.7(1.35) 3.6(2.80) 28821.5(18491.99)
HMAS-2 20% 17.0(9.0) 24.0(20.0) 355479.50(291054.50)
2x4 LLaMAC 90% 7.44(2.95) 10.56(7.54) 94029.44(68090.07)
HMAS-2 0% - - -
4x8 LLaMAC 90% 8.44(1.57) 12.11(2.51) 119773.55(32748.69)
@@
We conduct a comparative analysis between our method and the state-of-the-art solution, HMAS-2 Chen et al. (2023). For each scenario, we conduct tests on grid configurations of $2\times 2$, $2\times 4$, and $4\times 8$, respectively. Table 1 and Table 2 present a comprehensive performance comparison between the two methods, clearly demonstrating the overall superiority of our approach. In complex scenarios involving long-term iterative decision-making, LLaMAC exhibits a significantly higher success rate compared to HMAS-2. Furthermore, LLaMAC consistently achieves task completion in fewer interaction steps, highlighting the performance advantages of its employed strategies. Additionally, the TripletCritic facilitates the generation of superior initial suggestions, thereby reducing the need for feedback iterations and greatly enhancing token utilization efficiency.

### 09 Dec 2023 v2

@@ 1 Introduction
Relying on training from massive datasets to capture extensive common knowledge and having demonstrated certain reasoning capabilities, Large Language Models (LLMs) have been widely applied and explored across various domains, rapidly emerging as powerful tools Brown et al. (2020); Kojima et al. (2022); Ruan et al. (2023a); Yang et al. (2023). Examples of such applications include question-answering systems Mallen et al. (2023), common-sense reasoning Hao et al. (2023), programming Tian et al. (2023), and embodied intelligence Driess et al. (2023). Recently, in the fields of natural language processing (NLP) and multi-agent systems (MAS), numerous research endeavors are dedicated to exploring the collaborative accomplishment of tasks through the utilization of multiple LLM-based agents. These efforts leverage the methodologies of role-playing Li et al. (2023) and debate Chan et al. (2023) as means to facilitate synergy and effective coordination among the agents involved. However, the majority of existing works primarily concentrate on coordination among a limited number of agents. The application of LLMs for effective coordination in large-scale agent scenarios has been rarely discussed.
@@
Table 1: Comprehensive comparison of multi LLM-based agets methods. All approaches rely on either multi-agent debate or role-playing to accomplish decision-making tasks and solve NLP problems (task solver), or simulate collective behavior (community simulator).
Type Method Target Role Agents Num.
Muti-Agent Debate Debate (Du et al.) Task Solver 2 debaters 2
MAD (Liang et al.) 1 judge + 2 debaters 3
ChatEval (Chan et al.) multi debaters 5
Role Playing CAMEL (Li et al.) Task Solver 1 assistant + 1 user 2
AgentVerse (Chen et al.) 1 role assigner + 2-4 experts + 1 evaluater 6
Proagent (Zhang et al.) 2 cooks 2
LLaMAC (ours) 3 critic + 1-50 actors 50
Generative Agents (Park et al.) Community Simulator 25 agents 25
Werewolf Agents (Xu et al.) 7 players 7
ReCon (Wang et al.) 6 players 6
To this end, we present LLaMAC, a novel framework for achieving a comprehensive decision-making process in collaborative tasks involving large-scale LLM-based agents, drawing inspiration from the classical actor-critic reinforcement learning (RL) approach Konda and Tsitsiklis (1999). Within LLaMAC, we design a centralized critic which takes on the role of a coordinator, making suggestions to each actor based on their decision memory. Subsequently, the actors interact with the environment, receive assigned tasks, and execute corresponding actions. (1) To tackle the exploration-exploitation trade-off inherent in the decision-making process, we introduce a TripletCritic structure inspired by the distributional code for value in the brain Dabney et al. (2020). This structure coordinates multiple critics with the same objective but different preferences. (2) We also establish a comprehensive feedback mechanism that incorporates both internal feedback within the TripletCritic and external feedback between the LLM-based actors (i.e., agents) and the TripletCritic. This mechanism aims to mitigate hallucination issues and bolster the robustness of LLM decision-making. (3) To construct a token-efficient planning framework, we incorporate mechanisms such as redundant memory information deletion and external feedback judgment. These mechanisms effectively reduce the length of context tokens and minimize the frequency of LLM accesses.
We first evaluate the performance of our method on a system resource allocation task to demonstrate its ability to strike a balance between exploration and exploitation, as well as its capability in large-scale multi-agent decision-making tasks. We further deploy our method in a more complex robot grid transportation scenario to validate its planning and decision-making capabilities. Experimental results demonstrate that our method outperforms existing approaches in terms of final performance, token utilization efficiency, and policy stability. To the best of our knowledge, we are the first to apply LLMs to large-scale multi-agent decision-making tasks.
@@ 2 Related Work
2.1 Multi-Agent Cooperation
@@
2.2 Planning and Reasoning with LLM
Learning in massive corpora gives LLMs certain commonsense reasoning capabilities Kojima et al. (2022). Although there are still challenges in solving complex decision tasks, a large amount of work has proven that their methods can effectively improve the planning ability of LLMs Zelikman et al. (2022); Creswell et al. (2022). One line of research focuses on decomposing complex queries into sequential intermediate steps, known as Chain-of-Thought (CoT) Wei et al. (2022), to achieve accurate solutions. Another direction involves incorporating feedback mechanisms, showcasing their extensive capabilities in tackling complex decision-making challenges Wang et al. (2023b). Moreover, recent studies have begun to address this issue employing multiple LLMs. These approached are enhanced in their planning capabilities through techniques such as debate Chan et al. (2023); Liang et al. (2023) or role-playing Li et al. (2023); Hong et al. (2023). In the domain of decision-making, a subset of research utilizes prompting techniques to construct comprehensive processes covering perception, planning, and action, in cluding video games Zhang et al. (2023c), robot contro l Zhang et al. (2023d); Mandi et al. (2023), and open-world tasks Zhu et al. (2023); Gong et al. (2023). There are some studies about task planning and external tool usage Ruan et al. (2023b); Kong et al. (2023). However, as shown in Table 1, nearly all existing studies have only focused on tasks involving a small number of agents, with the involvement of a larger number of agents only observed in methods analyzing community simulations that do not require task-solving. However, it is worth noting that the existing studies, as outlined in Table 1, have predominantly concentrated on tasks involving a limited number of agents. The involvement of a larger number of agents has primarily been observed in methods that analyze community simulations, wherein task-solving is not a requirement. In light of this observation, our work uniquely emphasizes the application of language models in the realm of decision-making within large-scale multi-agent systems.
@@ 4.1.2 Results
[…]
HMAS-2 100% 9.9(2.74) 3.3(2.05) 49.9(17.98)
2x2 LLaMAC 100% 7.0(1.79) 2.0(1.26) 23.9(8.38)
HMAS-2 80% 15.5(6.09) 12.3(5.83) 158.4(107.84)
2x4 LLaMAC 100% 7.6(1.36) 4.3(1.42) 38.0(10.57)
HMAS-2 60% 30.6(9.70) 26.1(13.59) 599.3(245.40)
4x8 LLaMAC 100% 12.9(2.70) 10.7(3.35) 122.6(30.55)
[…]
HMAS-2 80% 7.0(5.0) 6.0(9.74) 76.1(116.66)
2x2 LLaMAC 100% 4.7(1.35) 3.6(2.80) 28.8(18.49)
HMAS-2 20% 17.0(9.0) 24.0(20.0) 355.5(291.05)
2x4 LLaMAC 90% 7.44(2.95) 10.56(7.54) 94.0(68.09)
HMAS-2 0% - - -
4x8 LLaMAC 90% 8.44(1.57) 12.11(2.51) 119.8(32.75)
@@ 4.2 Grid Transportation
We conduct a comparative analysis between our method and the state-of-the-art solution, HMAS-2 Chen et al. (2023b). For each scenario, we conduct tests on grid configurations of $2\times 2$, $2\times 4$, and $4\times 8$, respectively. Table 2 and Table 3 present a comprehensive performance comparison between the two methods, clearly demonstrating the overall superiority of our approach. In complex scenarios involving long-term iterative decision-making, LLaMAC exhibits a significantly higher success rate compared to HMAS-2. Furthermore, LLaMAC consistently achieves task completion in fewer interaction steps, highlighting the performance advantages of its employed strategies. Additionally, the TripletCritic facilitates the generation of superior initial suggestions, thereby reducing the need for feedback iterations and greatly enhancing token utilization efficiency.